OpenStudy (anonymous):

Calculate the value of sin 15 .

OpenStudy (anonymous):

\[\cos2 \theta =1-2 \sin^{2} \theta\] so \[\cos2 \theta =1-2 \sin^{2} \theta\] \[\sin \theta =\sqrt{\frac{1-\cos 2 \theta}{2}}\] put theta=15